OpenStudy (anonymous):

Factor the ff : 125-y3 = x3y3-93 = b3-1000 = Thanks :)

OpenStudy (anonymous):

\[ a^3 - b^3 =(a-b)(a^2 + ab +b^2) \]

OpenStudy (anonymous):

\[ 125 - y^3 = 5^3 -y^3=(5-y)(25+ 5 y + y^2) \] Mimic this one to finish the other two.

OpenStudy (anonymous):

Can you do the other two?

OpenStudy (anonymous):

What about the last two numbers please :(

OpenStudy (anonymous):

Are you sure that the second one is \[x^3 y^3 - 93 \]

OpenStudy (anonymous):

\[ b^3 - 1000=b^3 - 10^3=(b-10)(b^2 +10b + 100) \]